Spectroscopic observations of compact solar flares

H.E. Mason & Giulio Del Zanna

Department of Applied Mathematics and Theoretical Physics

University of Cambridge

in collaboration with A. Berlicki & B. Schmieder Observatoire de Paris, Meudon

Introduction

• The SOHO/Coronal Diagnostic Spectrometer provides essential spectroscopic information to understand the characteristics and dynamics of solar coronal features.

• We give a description of the main observational features and what types of measurements (densities, temperatures, flows) are obtained.

• These provide important observational constraints for flare modeling.

• CDS observations of a compact M1 flare that occurred on October 22 2002, also observed from the ground in chromospheric lines and by RHESSI.

• The flare occurred in a region of strong and mixed magnetic polarity, was compact and lasted only a short time.

Papers – I

- Berlicki et al., A&A

- Del Zanna et al. (2004) submitted

Papers – II

• On general aspects of quiescent ARs:

-Del Zanna & Mason 2003, A&A, 406, 1089 -Del Zanna 2003, A&A, 406, L5

• On chromospheric evaporation:

-Mason et al. 1986, ApJ, 309,435 -Zarro et al., 1988, ApJ, 333, L99 -Del Zanna et al. (COSPAR 2000), 2002, Adv. Space Res., 30, 551

• Time dependent and non-equilibrium effects: -Bradshaw, Del Zanna & Mason, A&A, 2004, in press

Overview I

Figure 1: THEMIS/MSDP in Na I D_{1} (15:41–16:29 UT) and line-of-sight magnetic field deduced from I and V Stokes overlaid with Hα contours observed by VTT/MSDP at 15:30:01 UT.

December 2004: compact solar flares – p.5/25

Overview II

• THEMIS observed continuously between 15:41–17:26 UT.

• TRACE and VTT/MSDP observations were performed during the whole flare.

• VTT/MSDP observations in Hα 9 scans between 15:30 and 16:17 UT

• TRACE high cadence (40 s) in the 195 Å band between 15:29 and 15:41 UT

Overview III - GOES

December 2004: compact solar flares – p.7/25

Overview III - GOES

December 2004: compact solar flares – p.8/25

The AR before the M1 flare

TRACE 195 Å with MDI contours

The M1 flare - TRACE (area Z1)

December 2004: compact solar flares – p.10/25

The M1 flare

CDS FOV - TRACE 195 Å + contours of Fe XIX intensity

The M1 flare - CDS

December 2004: compact solar flares – p.12/25

The M1 flare - CDS – Zoom in

December 2004: compact solar flares – p.13/25

Wavelength calibration

The wavelength calibration was remarkably stable. Uncertainty in the derivation of velocities of the order of 5–10 km/s.

Si XII  He I  Ne IV  Al XI  O IV  Ca X  Ne VI  Fe XII  Mg X  O V
λ_{st}  520.665 522.210  542.07  550.06  555.263 557.765 562.803  592.60  624.941 629.732
λ_{1}  520.675 522.208 542.031 550.070 555.252 557.758 562.791 592.590 624.936 629.735
v_{1} 5 -1 -21 5 -6 -3 -6 -5 -2 1
λ_{2}  520.664 522.211 542.061 550.066 555.260 557.769 562.812 592.597 624.932 629.747
v_{2}  0  0  -4  3  -1  1  5  -1  -4  6

Selected areas

Si XII velocity map, with the areas selected for further analysis. Areas A,B,C correspond to regions of strong blue-shifts, while area D of red-shift. Region LP corresponds to the area of peak Fe XIX intensity, while region L1 is where the central part of the flare loop is located. Region BG is a ‘background’ reference region.

December 2004: compact solar flares – p.15/25

The Fe XIX case

Our calibration, based on this small sample, provides a wavelength of 592.16 Å, in excellent agreement with Skylab measurements. Lawson & Peackok (1980) reviewed n = 2 → n = 2 transitions, and suggested an energy for the 2s 2 2p 4^{1}D_{2} level of 168770 cm^{−1}. → λ = 592.522 Å → the considerable difference of 183. km/s compared to our result.

Densities from O IV

December 2004: compact solar flares – p.17/25 December 2004: compact solar flares – p.18/25

Velocities, densities

Region  A  B  C  D  L1  LP  BG
He I v  13  -27  -30  -17  -17  -36  -22
I  2393  770  112  377  751  1293  16
F W HM  0.61  0.53 0.51  0.56  0.51  0.51  0.50
O IV(a) v  (33)  -21  -24  13  0  -20  0
I  (9818)  598  80  237  252  686  33
O IV(b) I  2333  111  -  -  -  -  -
N_{e} (O IV)  2 10^{11}  0.9 10^{11}  -  -  -  -  -
O V v  (23)  -25  -21  0  -12  -28  -4
I  (33226)  5038  721  2490  2840  6782  387
F W HM  (0.70)  0.50 0.50  0.55  0.50  0.50  0.50
Mg X v  (14)  -62  -51  29  -17  -37  -5
I  (5910)  1630  656  1205  1781  17093  118
F W HM  (0.80)  0.59 0.53  0.58  0.58  0.52  0.52
Si XII v  -64  -84  -54  17  -6  -40  -10
I  2756  1847  839  1378  2818  16748  63
F W HM  0.59  0.55 0.50  0.55  0.51  0.50  0.50
Fe XIX v  -141  -140  -56  22  -13  -26  -
I  294  272  694  1271  2721  7544  ≤ 0.5
F W HM  0.68  0.65 0.56  0.59  0.57  0.55  -

December 2004: compact solar flares – p.19/25

VTT Hα

Area log τ_{m} V_{0}[km s^{−1}]
1  −0.4 downflow (?)
2  −0.2 downflow (?)
3  +0.4  +10.0
4  +1.0  −1.9
5  +0.1  −3.8
6  −0.8  −9.4
7  −0.1  −5.6
8  0.0  −6.9

December 2004: compact solar flares – p.20/25

Later flare - GOES

December 2004: compact solar flares – p.21/25

Later flare - RHESSI

RHESSI counts rates (the 12–25 keV band are multiplied by a factor of 10).

Post-flare CDS raster 15:43–17:07 UT

December 2004: compact solar flares – p.23/25

Post-flare CDS raster – II

December 2004: compact solar flares – p.24/25

Summary of results

• Strong blue-shifts in lines emitted in the 1-8 MK range are observed at the bases of hot Fe XIX structures.

• The hot structures are also clearly visible in the TRACE 195 Å band • The hot structures match the thermal emission as seen by RHESSI • The peak Fe XIX emission, recorded between 15:35 and 15:43 UT,

occurred just above the region of strong mixed polarity.

• The TRACE 195 Å emission is co-spatial with the CDS Fe XIX emission, which suggests that the bulk of the TRACE 195 Å emission at the flare site is due to Fe XXIV.

• Transition region densities can be estimated using O IV.

Acknowledgements Support from PPARC is acknowledged. SOHO is a project of international collaboration between ESA and NASA. We acknowledge the use of the SURF, TRACE on-line databases.